In corn, a dihybrid for the recessive genes a and b is test-crossed. The distribution of the phenotypes is as follows:

A B 122
A b 118
a B 81
a b 79

a) What appears to be the distance between genes a and b?

b) Test your hypothesis with a chi-squared test. Do the genes appear to be sorting independently?
Answer #1

Answer:

a).

Expected testcross progeny would be 1:1:1:1 ratio according to Mendel experiments. But the given numbers of the progeny are not similar to 1:1:1:1 ratio. So, the genes are not in assorted independently and they are linked. Each expected progeny = 1/4

Recombination frequency (RF) = (no. of recombinant progeny / Total progeny)100

RF = (81+79/400)100 = 40%

RF(%) = Distance between the genes (m.u.)

The distance between the two genes = 40 m.u.

b).

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

AB

122

100.00

22.000

484.000

4.840

Ab

118

100.00

18.000

324.000

3.240

aB

81

100.00

-19.000

361.000

3.610

ab

79

100.00

-21.000

441.000

4.410

Total

400

400

16.100

Chi-square value = 16.10

Degrees of freedom = Number of phenotypes – 1

Df = 4-1=3

Critical value = 7.81

The chi-square value of 16.10 is greater than the critical value of 7.81. We can reject null hypothesis and the two genes are not assorted independently.